OpenStudy (anonymous):

A separate blueprint for the ramp uses a scale of 1 cm : 25 cm. The height of the ramp on the blueprint is ____ cm. (Enter only the number.)

OpenStudy (anonymous):

A separate blueprint for the ramp uses a scale of 1 cm : 25 cm. The height of the ramp on the blueprint is ____ cm. (Enter only the number.)

OpenStudy (anonymous):

Do you know how to write the proportion for 1/25 = ?

OpenStudy (anonymous):

no can you help??

OpenStudy (anonymous):

Notice that the height is 50 cm in the picture; therefore, \[\frac {1}{25} = \frac {50}{x}\]

OpenStudy (anonymous):

Now we can solve for x ...

OpenStudy (anonymous):

1x=1250?

OpenStudy (anonymous):

Would my answer be 1250?